OpenStudy (anonymous):

Can someone please help me!? In ΔABC, a = 7.2cm, b = 9.3cm, ∠A = 35° (6) a. Determine the number of possible solutions. b. Determine the measure of ∠B if it exists.

OpenStudy (anonymous):

i) Applying sine law of triangles, b/sin(<B) = a/sin(<A) ==> 9.3/sin(B) = 7.2/sin(35°) ii) Solving the above, sin(B) = 0.7409 This give <B = 47°48'; as well sin(180 - B) = sin(B); So another value for <B = 180° - 47°48' = 132°12' Thus <B = Either 47°48' or 132°12' Both values of <B are possible.

OpenStudy (anonymous):

i) Applying sine law of triangles, b/sin(<B) = a/sin(<A) ==> 9.3/sin(B) = 7.2/sin(35°) ii) Solving the above, sin(B) = 0.7409 This give <B = 47°48'; as well sin(180 - B) = sin(B); So another value for <B = 180° - 47°48' = 132°12' Thus <B = Either 47°48' or 132°12' Both values of <B are possible.